prove tan3A.tan2A.tanA = tan3A - tan2A - tanA


prove


tan3A.tan2A.tanA = tan3A - tan2A - tanA


Grade:11

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

tan(3A) = tan(2A+A)
tan(3A) = [tan(2A) + tan(A)] / [1 - tan(A)*tan(2A)]

cross multiply
tan(3A) * [1 - tan(A)*tan(2A)] = [tan(2A) + tan(A)]
tan(3A) - tan(A)*tan(2A)*tan(3A) = tan(2A) + tan(A)
tan(A)*tan(2A)*tan(3A) = tan(3A) - tan(2A) - tan(A)

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

vikas askiitian expert
13 years ago

tan2A can be written as tan(3A-A)

 

now we have tan(a-b)=tana-tanb/1+tanatanb

 

here a =2A & b=A

tan2A=tan(3A-A)=tan3A-tanA/1+tan3AtanA

 

tan2A(1+tan3Atan2A) = tan3A-tanA

 

tan2A + tan2Atan3AtanA = tan3A -tanA or

 

tan2Atan3AtanA = tan3A -tanA-tan2A

hence proved